# Learning from Heterogeneity: A Dynamic Learning Framework for Hypergraphs

### v1

@@ Abstract
Graph neural network (GNN) has gained increasing popularity in recent years owing to its capability and flexibility in modeling complex graph structure data. Among all graph learning methods, hypergraph learning is a technique for exploring the implicit higher-order correlations when training the embedding space of the graph. In this paper, we propose a hypergraph learning framework named Learning from Heterogeneity (LFH) that is capable of dynamic hyperedge construction and attentive embedding update utilizing the heterogeneity attributes of the graph. Specifically, in our framework, the high-quality features are first generated by the pairwise fusion strategy that utilizes explicit graph structure information when generating initial node embedding. Afterwards, a hypergraph is constructed through the dynamic grouping of implicit hyperedges, followed by the type-specific hypergraph learning process. To evaluate the effectiveness of our proposed framework, we conduct comprehensive experiments on several popular datasets with eleven state-of-the-art models on both node classification and link prediction tasks, which fall into categories of homogeneous pairwise graph learning, heterogeneous pairwise graph learning, and hypergraph learning. The experiment results demonstrate a significant performance gain (average 12.5% in node classification and 13.3% in link prediction) compared with recent state-of-the-art methods.
@@
Heterogeneous Hypergraph, Representation Learning, Hypergraph Generation, Classification.
@@ I Introduction
There are both standard graphs and hypergraphs. Fig. 1 illustrates (a) a standard graph where one edge connects two vertices regardless of the edge type and (b) a hypergraph with three hyperedges. Standard graphs can be modeled using many existing GNN models such as graph convolution network (GCN) [5], graph attention network (GAT) [6], and heterogeneous graph learning (PC-HGN) [7]. Most of these models aim to model the pairwise relationships in well-structured graph data. However, this may not work for hypergraphs, as a hypergraph can encompass any number of vertices (nodes) at each edge type, presenting higher-order correlations with rich underlying information. Pairwise graph learning models, in this case, become intractable. Specifically, hypergraph presents more complex non-pairwise relationships, where these implicit relationships are not only dyadic (pairwise) but rather triadic or even higher [8]. In general, hypergraph learning can be regarded as a generalized form of traditional graph learning.
Like pairwise graph learning models like GCN, hypergraph learning can also be seen as a process of uncovering and passing the structure information among different hyperedges, during which low-dimensional node/graph level embedding representations can be learned and used for downstream tasks, such as node classification [9], link prediction [10], and clustering [11].
Figure 1: Comparison between (a) a standard graph with 5 nodes and 4 edges, and (b) a hypergraph with 3 hyperedges (the three colored node groups).
@@
Owing to the great potential of exploring high-order correlations among the data, hypergraph learning has drawn increasing attention from the academic community and has been employed in applications such as image classification [12], video segmentation [13], and hyperspectral image analysis [14]. How to construct hyperedges lies at the core of hypergraph learning. Existing works generally take two approaches, i.e., attribute-based construction and neighborhood-based construction. The early work focused on constructing the hypergraph based on the features of each vertex [15]. This strategy instructs the model to use only the feature information which inevitably sacrifices certain generality from the topological perspective. The neighborhood-based strategy addresses this weakness by enabling the locality concept. Specifically, a hyperedge is constructed through clustering a master vertex and its neighbor vertices [16]. However, the clustering method (e.g., k-nearest neighbors (k-NN)) separates hypergraph construction from the graph learning process, and thus is sensitive to noisy data, especially in visual classification tasks. Moreover, most neighborhood-based models fail to incorporate topological heterogeneity into the learning process, which limits the learning to a single type of hyperedges.
@@
  • We design a dynamic learning process to generate different types of implicit hyperedges to construct the hypergraph and dynamically adapt the hypergraph construction during the learning process. The embedding is updated iteratively through a type-specific multi-head attention mechanism, so as to encode the heterogeneous attributes into the embedding space.
@@
The remainder of this paper is organized as follows. Section 2 reviews the related works of graph learning on homogenous, heterogenous, and hypergraphs. Section 3 provides the preliminary knowledge of hypergraphs, while Section 4 introduces the detail of our proposed LFH framework, along with analyses and discussions. The experimental results are reported and analyzed in Section 5. Finally, Section 6 concludes the paper.
@@ II-C Hypergraph Learning
Hypergraph learning explores the high-order correlations in the data, which extends the traditional graph learning models to a high dimensional and more complete nonlinear space. It offers a promising solution for analyzing complex structured data with satisfactory performance in practice [16]. The construction of the hypergraph from the given data is a key step for hypergraph learning, which significantly affects the final learning performance.
@@
Zhang et al. proposed dynamic hypergraph structure learning (DHSL) to update hypergraph structure iteratively during the learning process [27]. It is essential to make dynamic modifications to the initial hypergraph structures from adjusted feature embedding. Feng et al. presented a framework named hypergraph neural network (HGNN) for handling complex and high-order correlations. In HGNN, the complex data correlation is formulated in a hypergraph structure, and a hyperedge convolution operation is used to exploit the high-order data correlation for representation learning [9].
@@
To exploit high-order relations among the features, Jiang et al. proposed a dynamic hypergraph neural networks (DHGNN) framework that is composed of dynamic hypergraph construction (DHG) and hypergraph convolution (HGC). DHG utilizes the k-NN method to generate the basic hyperedge and extends the adjacent hyperedge set by k-means clustering, with which the local and global relations can be extracted. HGC is designed to encode high-order data relations in the hypergraph structure [28]. Cai et al. introduced a hypergraph structure learning (HSL) framework to optimize the hypergraph structure and HGNNs simultaneously in an end-to-end way. To efficiently learn the hypergraph structure, HSL adopts a hyperedge sampling strategy to prune the redundant hyperedges, which is followed by an incident node sampling for pruning irrelevant incident nodes and discovering potential implicit connections. The consistency between the optimized structure and the original structure is maintained by the intra-hyperedge contrastive learning module [29]. Gao et al. proposed a tensor-based dynamic hypergraph learning (t-DHL) model to efficiently learn dynamic hypergraphs. t-DHL utilizes a tensor representation to characterize the dynamic hypergraph structure more flexibly. During the optimization of the tensor representation, not only the weights but also the number and order of hyperedges can be adjusted. As an extended version of HGNN, a general hypergraph neural network framework named HGNN+ was proposed for modeling high-order representation among the data, which is achieved by bridging multi-modal/multi-type data and hyperedge with hyperedge groups [30].
@@
Recent works mostly focus on neighborhood-based construction strategies when constructing hyperedges in the hypergraph. Specifically, hyperedge is often constructed through clustering a master vertex and its neighbor vertices. However, for the commonly adopted clustering methods like k-NN [12, 28, 9], they treat hypergraph construction and graph learning separately. As a result, these models are susceptible to noisy data which limits their applications. Apart from that, most neighborhood-based models fail to incorporate the heterogeneity of the topology, leading to a single type hyperedge [16]. In this work, we propose a heterogeneous hypergraph learning framework LFH to solve the above mentioned issues. LFH constructs the hyperedge based on different edge types, which can then be integrated into the learning process to improve the learned representations.
@@ III Preliminaries and problem formulation
Figure 2: An illustrative example of the ACM dataset: (a) Heterogeneous pairwise graph including four node types (Author, Paper, Subject) and four edge types (Write, Written by, Belong to, Contain). (b) A heterogeneous hypergraph that models two implicit data relations: preference for subjects and the relation among the subjects.
@@
Figure 2 provides an illustrative comparison between a heterogeneous pairwise graph and a hypergraph. Figure 2 (a) sketches the ACM dataset composed of three node types (Author, Paper, Subject) and four edge types (Write, Written by, Belong to, Contain). By contrast, a heterogeneous hypergraph is capable of modeling a more complex graph structure by constructing different types of hyperedges, in which every hyperedge can connect any number of nodes to model a specific type of data relationship. As shown in Figure 2 (b), the purple hyperedge models an author’s preference for subjects, while the yellow one models the relation among different subjects.
@@
TABLE I: Table of key notations
@@ IV-B Heterogeneous Hyperedge Generation
The special structure of hyperedges enables hypergraphs to encode high-order data relations. By encircling a specific set of nodes with common attributes or close relationships within one hyperedge, we can effectively represent local group information among the nodes. Therefore, it is essential to discover the related nodes for different hyperedges. To facilitate hyperedge construction, we define two types of nodes: master node and slave node. A master node $v^{m}_{i}\in\mathcal{V}$ serves as an anchor when constructing a hyperedge, which dynamically encircles a set of slave nodes $\{v^{s}_{i}\}$ to jointly form the hyperedge. The collection of participant nodes is denoted as $S\left(v^{m}_{i}\right)$, which is united with master node $v_{i}^{m}$ to represent a complete data relation within a hyperedge. Previous works have adopted the $k$-NN method to select the k-th nearest nodes to form a hyperedge [27, 31]. Alternative approaches solve an independent optimization problem to generate the hyperedges [12, 32], however, this solution fails to incorporate the heterogeneity of the hyperedges, and thus cannot be integrated into the representation learning process.
@@
To address the above challenges, here we propose a dynamic learning process that generates heterogeneous hyperedges and adaptively reconstructs the hyperedge in different types. Specifically, for each master node $v^{m}_{i}\in\mathcal{V}$, our framework encloses the related slave nodes of various types to encode heterogeneous topological structures with implicit relationships, based on different attributes among the nodes. Suppose the master node $v^{m}_{i}$ has a set of $n_{i}$ candidate slave nodes $\tilde{S}^{k}\left(v^{m}_{i}\right)$, where $k=\left[1,...,n_{i}\right]$ with different types. The hyperedge can be reconstructed based on the node types of the master node and candidate slave node sets. The type of the candidate slave node set can be defined as $\mathcal{T}\left(\tilde{S}^{k}\left(v^{m}_{i}\right)\right)$ and clearly is a subset of node type $\mathcal{T}_{v}$. Each candidate slave node set $\tilde{S}^{k}\left(v^{m}_{i}\right)$ contains all nodes with one or several node types, which is denoted as $\tilde{S}^{k}\left(v^{m}_{i}\right)=\left\{v|\phi\left(v\right)\in\mathcal{T}\left(\tilde{S}^{k}\left(v^{m}_{i}\right)\right),v\neq v_{i}^{m}\right\}$. Since each hyperedge is generated from the corresponding candidate slave node set of a certain type, the type number of the hyperedge is equal to the type number of the candidate slave node set. An exemplary hyperedge construction is demonstrated in the middle part of Figure 3, where the master node encompasses $|\mathcal{T}_{v}|$ candidate slave node sets with distinct types. For each type, a hyperedge (represented as the coloured triangle) is dynamically constructed. Mathematically, this process of generating a hyperedge from a master node $v_{i}^{m}$ and one of its candidate slave node set $\tilde{S}^{k}\left({v_{i}^{m}}\right)$ can be denoted as:
@@
Figure 4: Type-specific multi-head attention on node embedding update
@@ IV-C2 Multi-head Attention Node Embedding Updating
where the dimension of $\emph{X}\left(v^{m}_{i}\right)$ is $\mathcal{R}^{d\times 1}$, and $Q\text{-}Proj^{h}_{\phi\left(v^{m}_{i}\right)}\in\mathcal{R}^{\frac{d}{K}\times d}$ represents the projection matrix on the $h$-th attention head, $h\in\left[1,K\right]$. $K$ is the number of attention heads. Note the projection matrices are distinguished on hyperedge types. Similarly, we also project the hyperedge $e^{1}\left(v^{m}_{i}\right)$ through $K\text{-}Proj^{h}_{\psi\left(e^{1}\left(v^{m}_{i}\right)\right)}$ into $\emph{K}^{h}\left(e^{1}\left(v^{m}_{i}\right)\right)$ as:
@@
[…]

### v2

@@ Abstract
Graph neural network (GNN) has gained increasing popularity in recent years owing to its capability and flexibility in modeling complex graph structure data. Among all graph learning methods, hypergraph learning is a technique for exploring the implicit higher-order correlations when training the embedding space of the graph. In this paper, we propose a hypergraph learning framework named Learning from Heterogeneity (LFH) that is capable of dynamic hyperedge construction and attentive embedding update utilizing the heterogeneity attributes of the graph. Specifically, in our framework, the high-quality features are first generated by the pairwise fusion strategy that utilizes explicit graph structure information when generating initial node embedding. Afterwards, a hypergraph is constructed through the dynamic grouping of implicit hyperedges, followed by the type-specific hypergraph learning process. To evaluate the effectiveness of our proposed framework, we conduct comprehensive experiments on several popular datasets with twelve state-of-the-art models on both node classification and link prediction tasks, which fall into categories of homogeneous pairwise graph learning, heterogeneous pairwise graph learning, and hypergraph learning. The experiment results demonstrate a significant performance gain (an average of 12.9% in node classification and 12.8% in link prediction) compared with recent state-of-the-art methods.
@@
Heterogeneous Hypergraph, Representation Learning, Hypergraph Generation, Classification
@@ I Introduction
Despite the exciting progress made by traditional GNNs, such as GCN [5], GAT [6], and GraphSage [7], in learning a low-dimensional latent representation of pairwise graphs, effectively dealing with heterogeneity and capturing high-order data correlations remain critical challenges for graph learning. Taking a bibliographic network as an example, Fig. 1 (a) sketches the ACM dataset composed of three node types (Author, Paper, Subject) and four edge types (Write, Written by, Belong to, Contain). The edges in the heterogeneous pairwise graph only depict disparate relations between two nodes. These relations are often explicit and pre-defined within the dataset. For instance, the Written relation reveals the paper’s authorship, and the Belong to relation indicates the paper’s research field. The heterogeneous pairwise graph can be modeled by most of the aforementioned traditional GNNs and heterogeneous graph learning methods, such as PC-HGN [8], HAN [9] and HGT [10]. However, such pairwise relations may fail to uncover underlying high-order relations, such as which subjects are more relevant to an author. Pairwise graph learning methods, in this case, become intractable. By contrast, hypergraph presents more complex non-pairwise relationships, where these implicit relationships are not only dyadic (pairwise) but rather triadic or even higher [11]. Specifically, a heterogeneous hypergraph is capable of modeling a more complex graph structure by constructing different types of hyperedges, in which every hyperedge can connect any number of nodes to model a specific type of data relationship. Thus, the hyperedges in the heterogeneous hypergraph are able to capture the implicit relations by clustering different groups of nodes, from which high-order information can be better encoded. As shown in Fig. 1 (b), the purple hyperedge models an author’s preference for subjects, while the yellow one models the relation among different subjects. In general, hypergraph learning can be regarded as a generalized form of traditional graph learning.
Figure 1: An illustrative example of the ACM dataset: (a) Heterogeneous pairwise graph including three node types (Author, Paper, Subject) and four edge types (Write, Written by, Belong to, Contain). (b) A heterogeneous hypergraph that models two implicit data relations: preference for subjects and the relation among the subjects.
@@
Owing to the great potential of exploring high-order correlations among the data, hypergraph learning has drawn increasing attention from the academic community and has been employed in applications such as image classification [12], video segmentation [13], and hyperspectral image analysis [14]. How to construct hyperedges lies at the core of hypergraph learning. Solutions to hypergraph construction in the existing works are generally divided into two categories, i.e., attribute-based construction and neighborhood-based construction. The early work focused on constructing the hypergraph based on the features of each vertex [15]. Recently, heterogeneous-hypergraph-based models (HeteHG-VAE [16], Meta-HGT [17]) utilize pre-defined relations, such as multiple authors co-authoring a paper in citation dataset, to construct the heterogeneous hypergraph. This strategy instructs the model to use only the feature information which inevitably sacrifices certain generality from the topological perspective, and makes the model inapplicable in other datasets. The neighborhood-based strategy addresses this weakness by enabling the locality concept. Specifically, a hyperedge is constructed through clustering a master vertex and its neighbor vertices [18]. However, the clustering method (e.g., k-nearest neighbors (k-NN) [19]) separates hypergraph construction from the graph learning process, and thus is sensitive to noisy data, especially in visual classification tasks. Moreover, most neighborhood-based models fail to incorporate topological heterogeneity into the learning process, which limits the learning to a single type of hyperedges. In order to overcome this, general hypergraph neural network (HGNN+) [20] designs four types of hyperedge generation approaches, including pairwise edge-based, attribute-based, k-hop neighbors-based and feature-based, to form the hypergraph. Nevertheless, HGNN+ separates the hypergraph construction from the graph learning process and only uses a trainable vector to represent the weight of all hyperedges (i.e., a weighting scalar for each hyperedge), without considering the heterogeneity information. Heterogeneity modelling is another challenge in hypergraph learning [21]. Existing heterogeneous hypergraph models like HHGSA [22] and HWNN [23] treat the heterogeneous hypergraph as multiple homogeneous hypergraphs, without leveraging the cross-type interactions among diverse hyperedge and node types. Additionally, hypergraph learning (HL) [24] models the representation of hypergraph by formulating an optimization problem, which partitions nodes into several parts and constructs the objective function to minimize the supervised loss and the partition cost when generating hyperedges. Yet, it is proven that the optimization is NP-complete, leading to a high computational complexity [19].
@@
  • We design a dynamic learning process to generate different types of implicit hyperedges to construct the hypergraph and dynamically adapt the hypergraph construction during the learning process. The embedding is updated iteratively through a type-specific multi-head attention mechanism, which adaptively learns the importance of heterogeneous hyperedges during node embedding updates, thereby encoding the heterogeneity of the topology into the embedding space.
@@
  • We have made our source code publicly available to contribute further to advancements in this field. The source code is available at https://github.com/paperCodeEric/LFH.
The remainder of this paper is organized as follows. Section II reviews the related works of graph learning on homogenous, heterogenous, and hypergraphs. Section III provides the preliminary knowledge of hypergraphs, while Section IV introduces the detail of our proposed LFH framework, along with analyzes and discussions. The experimental results are reported and analyzed in Section V. Finally, Section VI concludes the paper.
@@ II-C Hypergraph Learning
Hypergraph learning explores the high-order correlations in the data, which extends the traditional graph learning models to a high dimensional and nonlinear space. For example, Sun et al. designed a hypergraph framework for human behavioral analysis to indicate the interactions between individuals and their environments. Since each edge in the hyperedges contains multiple nodes information, hypergraph learning is suitable to analyze rich sociological criteria effectively [32]. As a result, hypergraph learning offers a promising solution for analyzing complex structured data with satisfactory performance in practice [18]. The construction of the hypergraph from the given data is a key step for hypergraph learning, which significantly affects the final learning performance.
@@
Guo et al. designed a representation learning framework to enhance the learning of individuals’ preferences from their friends’ preferences for recommendation system. To exploit the group similarity for an accurate group representation, all groups are connected as hypergraph, and the task of group preference learning are treated as embedding hyperedges in a hypergraph, where an inductive hyperedge embedding method is introduced [33]. Zhang et al. proposed dynamic hypergraph structure learning (DHSL) to update hypergraph structure iteratively during the learning process [34]. It is essential to make dynamic modifications to the initial hypergraph structures from adjusted feature embedding. Feng et al. presented a framework named hypergraph neural network (HGNN) for handling complex and high-order correlations. In HGNN, the complex data correlation is formulated in a hypergraph structure, and a hyperedge convolution operation is used to exploit the high-order data correlation for representation learning [19].
@@
To exploit high-order relations among the features, Jiang et al. proposed a dynamic hypergraph neural networks (DHGNN) framework that is composed of dynamic hypergraph construction (DHG) and hypergraph convolution (HGC). DHG utilizes the k-NN method to generate the basic hyperedge and extends the adjacent hyperedge set by k-means clustering, with which the local and global relations can be extracted. HGC is designed to encode high-order data relations in the hypergraph structure [35]. Luo et al. proposed a hypergraph framework named SHINE, which can effectively learn subgraph representations by simultaneously utilizing the learned representations and inductively infer representations for subgraph predictions for genetic medicine [36]. Cai et al. introduced a hypergraph structure learning (HSL) framework to optimize the hypergraph structure and HGNNs simultaneously in an end-to-end way. To efficiently learn the hypergraph structure, HSL adopts a hyperedge sampling strategy to prune the redundant hyperedges, which is followed by an incident node sampling for pruning irrelevant incident nodes and discovering potential implicit connections. The consistency between the optimized structure and the original structure is maintained by the intra-hyperedge contrastive learning module [37]. Gao et al. proposed a tensor-based dynamic hypergraph learning (t-DHL) model to efficiently learn dynamic hypergraphs. t-DHL utilizes a tensor representation to characterize the dynamic hypergraph structure more flexibly. During the optimization of the tensor representation, not only the weights but also the number and order of hyperedges can be adjusted. As an extended version of HGNN, a general hypergraph neural network framework named HGNN+ was proposed for modeling high-order representation among the data, which is achieved by bridging multi-modal/multi-type data and hyperedge with hyperedge groups [20]. Liu et al. proposed a meta-path-aware hypergraph transformer (Meta-HGT) to encode heterogeneous information network embedding. Meta-HGT constructs the heterogeneous hypergraph based on pre-defined meta-paths, such as multiple authors co-authoring a paper, in the citation graph [17]. Fan et al. introduced a heterogeneous hypergraph variational autoencoder (HeteHG-VAE) to learn the latent topology information of the heterogeneous hypergraph, which is constructed from pre-defined high-order relations, specifically for the link prediction task. HeteHG-VAE designs a hyperedge attention module that employs a non-linear attention scheme. This scheme learns the importance for different types of nodes within each hyperedge during the process of encoding the hyperedge embedding [16]. Khan et al. proposed a heterogeneous hypergraph neural network for social recommendation using attention network (HHGSA), which treats the heterogeneous hypergraph as multiple homogeneous hypergraphs. HHGSA directly employs the hypergraph attention network (HGAT) [38] to encode node and hyperedge embeddings, and then obtains the final representations of users and items through simple aggregation [22].
@@
Recent works mostly focus on neighborhood-based construction strategies when constructing hyperedges in the hypergraph. Specifically, hyperedge is often constructed through clustering a master vertex and its neighbor vertices [39]. However, for the commonly adopted clustering methods like k-NN [12, 35, 19], they treat hypergraph construction and graph learning separately. As a result, these models are susceptible to noisy data which limits their applications. Apart from that, most neighborhood-based models fail to incorporate the heterogeneity of the topology, leading to a single type hyperedge [18]. In this work, we propose a heterogeneous hypergraph learning framework LFH to solve the above mentioned issues. LFH constructs the hyperedge based on different edge types, which can then be integrated into the learning process to improve the learned representations.
@@ III Preliminaries and problem formulation
TABLE I: Table of key notations.
@@ IV-B Heterogeneous Hyperedge Generation
The special structure of hyperedges enables hypergraphs to encode high-order data relations. By encircling a specific set of nodes with common attributes or close relationships within one hyperedge, we can effectively represent local group information among the nodes. Therefore, it is essential to discover the related nodes for different hyperedges. To facilitate hyperedge construction, we define two types of nodes: master node and slave node. A master node $v^{m}_{i}\in\mathcal{V}$ serves as an anchor when constructing a hyperedge, which dynamically encircles a set of slave nodes $\{v^{s}_{i}\}$ to jointly form the hyperedge. The collection of participant nodes is denoted as $S\left(v^{m}_{i}\right)$, which is united with master node $v_{i}^{m}$ to represent a complete data relation within a hyperedge. Previous works have adopted the $k$-NN method to select the k-th nearest nodes to form a hyperedge [34, 40]. Alternative approaches solve an independent optimization problem to generate the hyperedges [12, 41], however, this solution fails to incorporate the heterogeneity of the hyperedges into the integrated optimization process.
@@
To address the above challenges, here we propose a dynamic learning process that generates heterogeneous hyperedges and adaptively reconstructs the hyperedge in different types. Specifically, for each master node $v^{m}_{i}\in\mathcal{V}$, our framework encloses the related slave nodes of various types to encode heterogeneous topological structures with implicit relationships, based on different attributes among the nodes. Suppose the master node $v^{m}_{i}$ has a set of $n_{i}$ candidate slave nodes $\tilde{S}^{k}\left(v^{m}_{i}\right)$, where $k=\left[1,...,n_{i}\right]$ with different types. The hyperedge can be reconstructed based on the node types of the master node and candidate slave node sets. The type of the candidate slave node set can be defined as $\mathcal{T}\left(\tilde{S}^{k}\left(v^{m}_{i}\right)\right)$ and clearly is a subset of node type $\mathcal{T}_{v}$. Each candidate slave node set $\tilde{S}^{k}\left(v^{m}_{i}\right)$ contains all nodes with one or several node types, which is denoted as $\tilde{S}^{k}\left(v^{m}_{i}\right)=\left\{v|\phi\left(v\right)\in\mathcal{T}\left(\tilde{S}^{k}\left(v^{m}_{i}\right)\right),v\neq v_{i}^{m}\right\}$. Since each hyperedge is generated from the corresponding candidate slave node set of a certain type, the type number of the hyperedge is equal to the type number of the candidate slave node set. An exemplary hyperedge construction is demonstrated in the middle part of Figure 2, wherein the master node encompasses $|\mathcal{T}_{v}|$ candidate slave node sets with distinct types, each of which is composed of all nodes with one single type. For each type, a hyperedge (represented as the coloured triangle) is dynamically constructed. Mathematically, this process of generating a hyperedge from a master node $v_{i}^{m}$ and one of its candidate slave node set $\tilde{S}^{k}\left({v_{i}^{m}}\right)$ can be denoted as:
@@
Figure 3: Type-specific multi-head attention on node embedding update.
@@ IV-C2 Multi-head Attention Node Embedding Updating
where the dimension of $\emph{X}\left(v^{m}_{i}\right)$ is $\mathcal{R}^{d\times 1}$, and $Q\text{-}Proj^{h}_{\phi\left(v^{m}_{i}\right)}\in\mathcal{R}^{\frac{d}{K}\times d}$ represents the projection matrix on the $h$-th attention head, $h\in\left[1,K\right]$. $K$ is the number of attention heads. The master node is projected to generate the $h$-th query vector $\emph{Q}^{h}\left(v^{m}_{i}\right)$ by the projection matrix $Q\text{-}Proj^{h}_{\phi\left(v^{m}_{i}\right)}$. Note the projection matrices are distinguished on hyperedge types. Similarly, we also project the hyperedge $e^{1}\left(v^{m}_{i}\right)$ through $K\text{-}Proj^{h}_{\psi\left(e^{1}\left(v^{m}_{i}\right)\right)}$ into $\emph{K}^{h}\left(e^{1}\left(v^{m}_{i}\right)\right)$ as:
@@
$\emph{Z}^{\mathcal{V}}\left(v^{m}_{i}\right)=\textit{MLP}\left(\mathop{\operatorname*{\mathchoice{\Big{\|}}{\big{\|}}{\|}{\|}}}\limits_{h\in\left[1,K\right]}\emph{z}^{h}_{i}\right)$ (15)